the sum of two numbers is 23. the product of the two same numbers is 90. what are those two numbers
Found 2 solutions by nyc_function, edjones:
Answer by nyc_function(2741)

This is a system of linear equations in two variables.
Let x = one number
Let y = the other number

x + y = 23
(x)(y) = 90
Can you finish now?

Answer by edjones(8007)

x+y=23
x=23-y
.
xy=90
y(23-y)=90
23y-y^2=90
y^2-23y+90=0
(y-18)(y-5)=0
When y=18, x=5
When y=5, x=18
.
Ed